OpenStudy (anonymous):

Solve. 3/2x – 4 = 16

OpenStudy (jamesj):

Is this 3 divided by 2x or (3/2) times x ?

OpenStudy (anonymous):

(3/2) times x

OpenStudy (anonymous):

3/2(x) = 20 x = 20*(2/3) x = 40/3 = 13.333